### v1

# Retrieval-Oriented Knowledge for Click-Through Rate Prediction

Abstract.
Click-through rate (CTR) prediction plays an important role in personalized recommendations. Recently, sample-level retrieval-based models (e.g., RIM) have achieved remarkable performance by retrieving and aggregating relevant samples. However, their inefficiency at the inference stage makes them impractical for industrial applications. To overcome this issue, this paper proposes a universal plug-and-play retrieval-oriented knowledge (ROK) framework. Specifically, a knowledge base, consisting of a retrieval-oriented embedding layer and a knowledge encoder, is designed to preserve and imitate the retrieved & aggregated representations in a decomposition-reconstruction paradigm. Knowledge distillation and contrastive learning methods are utilized to optimize the knowledge base, and the learned retrieval-enhanced representations can be integrated with arbitrary CTR models in both instance-wise and feature-wise manners. Extensive experiments on three large-scale datasets show that ROK achieves competitive performance with the retrieval-based CTR models while reserving superior inference efficiency and model compatibility.
copyright: acmcopyrightjournalyear: 2023doi: XXXXXXX.XXXXXXXconference: xx; Preprint; price: xisbn: xxx
1. Introduction
Click-through rate (CTR) prediction is a key component of many personalized online services, such as recommender systems (Xi et al., 2023) and web search (Lin et al., 2021; Fu et al., 2023; Dai et al., 2021). It seeks to estimate the probability of a user’s click given a particular context (Zhang et al., 2021). The CTR models could be mainly classified into two categories. The first category is the feature interaction-based methods. The core idea of these methods is to capture the high-order feature interactions across multiple fields with different operators (e.g., product (Qu et al., 2018; Wang et al., 2021; Guo et al., 2017; Chen et al., 2021b), convolution (Xin et al., 2019; Liu et al., 2019), and attention (Song et al., 2019; Xiao et al., 2017)). The second category is the user behavior modeling methods that leverage different architectures (e.g., RNN (Hidasi and Karatzoglou, 2018; Hidasi et al., 2015), CNN (Tang and Wang, 2018), attention (Zhou et al., 2019, 2018), memory bank (Pi et al., 2019; Ren et al., 2019)) to extract informative knowledge from user behavior sequences for final CTR prediction.
To further enhance the performance, UBR4CTR (Qin et al., 2020) and SIM (Pi et al., 2020) retrieve useful behaviors from the user’s behavior history (i.e., clicked items), reducing the potential noise in user behavior sequences. Subsequent studies (Li et al., 2022b) have improved these methods’ efficiency using hashing functions (Chen et al., 2021a) and parallel retrieval execution (Cao et al., 2022) during inference. Recent studies (Qin et al., 2021; Du et al., 2022) further expand these retrieval-based methods from item-level retrieval to sample-level retrieval, applicable to general CTR prediction settings. Instead of retrieving similar items from the user history, RIM (Qin et al., 2021) adopts the idea of $k$ nearest neighbor ($k$NN) and designs a sample-centric retrieval method, which aggregates the relevant data samples retrieved from the search pool (e.g., the whole training dataset). PET (Du et al., 2022) constructs a hypergraph over the retrieved data instances and performs message propagation to get a better representation of the target data for final CTR prediction.
Although sample-level retrieval-based methods bring impressive performance enhancement, they have to perform instance-wise comparisons between the target data sample and each candidate sample in the search pool (usually million or even billion level). This leads to extreme inefficiency problems during inference, making it impractical for industrial applications, as illustrated in the left part of Figure 1. Recent work, such as DERT (Zheng et al., 2023), has improved the retrieval mechanisms by using vector retrieval techniques, enhancing efficiency. However, DERT’s dependency on the RIM Encoder and its $O(\log N)$ retrieval time complexity, along with the additional cost of encoding numerous samples in the retrieval pool, limits its scalability and effectiveness in large-scale applications.
This work introduces the Retrieval-Oriented Knowledge (ROK) framework to tackle the inference inefficiency problem of sample-level retrieval-based methods, with two stages: Retrieval-Oriented Knowledge Construction and Knowledge Utilization. Initially, we pre-train a sample-level retrieval-based model, such as RIM(Qin et al., 2021). To resolve the inefficient inference issue, ROK tries to learn a neural network-based Knowledge Base to imitate the aggregated representations from the pre-trained retrieval-based methods via a Retrieval Imitation module, whose simple diagram is shown in Figure 1. The Knowledge Base, built on a novel decomposition-reconstruction paradigm, where a Retrieval-Oriented Embedding Layer captures the feature-wise embedding and a Knowledge Encoder reconstructs the instance-wise aggregated representations (i.e., retrieval-enhanced representation). In this way, instead of the time-consuming retrieval, efficient inference of neural networks could be adopted to get the approximated aggregated representations. Additionally, we introduce a Contrastive Regularization module to ensure learning stability and prevent model collapse. At the Knowledge Utilization stage, ROK designs two approaches (i.e., instance-wise and feature-wise) to integrate the retrieval-enhanced representations with various CTR models, thus providing high inference efficiency.
The main contributions of this paper are as follows:
  • In this work, we introduce ROK, a pioneering framework that transforms sample-level retrieval-based methods —previously deemed infeasible in the industry into a practical solution. Our approach innovatively leverages a neural network-based Knowledge Base to efficiently store the retrieval-enhance representations and hence obviates the need for time-intensive retrieval & aggregation during inference.
  • We use knowledge distillation and contrastive learning to optimizie the knowledge base, allowing the integration of retrieval-enhanced representations with various CTR models, both at the instance and feature levels.
  • Extensive experiments conducted across three large-scale datasets have shown that ROK not only achieves competitive performance when compared to existing retrieval-based CTR methods but also reserves superior inference efficiency. Moreover, ROK enhances the performance of various CTR methods due to its exceptional compatibility. This success underscores the effectiveness of the neural knowledge model as a compact surrogate for the retrieval pool.
Figure 1. Comparison between traditional sample-level retrieval framework and ROK.
Figure 2. Overall framework of ROK. The process begins with the Retrieval-Oriented Knowledge Construction stage, where a sample-level retrieval-based model has already been pre-trained, and the retrieval and aggregation results are offline pre-stored. A tailored knowledge base is then designed to imitate the aggregation results, with Contrastive Regularization ensuring stabilization during the training phase. This design allows the time-consuming retrieval process to be substituted with the swift forward propagation of a neural network. In the subsequent Knowledge Utilization stage, the framework introduces the retrieval-enhanced representation into backbone CTR models in both an instance-wise and feature-wise approach, ensuring the maintenance of sample-level retrieval abilities while circumventing the complexities of retrieval and aggregation operations.
2. Related Work
2.1. CTR Prediction
For the click-through rate (CTR) prediction, models can essentially be divided into two main categories: those focusing on feature interaction and those centering on user behavior modeling. The first category is the feature interaction-based methods, evolving from foundational works such as POLY2 (Chang et al., 2010) and Factorization Machines (FM) (Rendle, 2010). With the integration of Deep Neural Networks (DNNs), a variety of sophisticated deep feature interaction models have been proposed. These models aim to capture high-order feature interactions across different fields by employing various operations, such as the product operation(Qu et al., 2018; Wang et al., 2021; Guo et al., 2017; Chen et al., 2021b; Li et al., 2022a), convolution (Xin et al., 2019; Liu et al., 2019), and attention mechanisms (Song et al., 2019; Xiao et al., 2017). The key innovation of these models lies in their ability to identify and utilize complex interactions between a multitude of features to enhance the predictive performance of CTR models.
Besides, user behavior modeling is another core technique for CTR prediction that mines user preferences from historical interaction behaviors meticulously. To better extract informative knowledge from user’s behavior sequence, various network structures have been utilized, including Recurrent Neural Networks (RNNs), Convolutional Neural Networks (CNNs), Attention Networks, and Memory Networks. GRU4Rec (Hidasi et al., 2015) designs Gated Recurrent Units (GRUs) to capture the preference-evolving relationship while Caser (Tang and Wang, 2018) leverages the horizontal and vertical convolution to model skip behaviors at both union-level and point-level. Moreover, the attention mechanism is the most popular method for modeling item dependencies, and several influential works have been proposed, including SASRec (Kang and McAuley, 2018), DIN (Zhou et al., 2018), DIEN (Zhou et al., 2019), and BERT4Rec (Sun et al., 2019). Among these, DIN and DIEN leverage the target-attention network to identify important historical items, while SASRec and BERT4Rec apply the self-attention network with Transformer architecture to excavate behavior dependencies. Besides, memory-based methods (Jiang et al., 2020) are also proposed to store user behavior representations in a read-write manner.
2.2. Retrieval-Augmented Recommendation
To further enhance the performance of CTR prediction, the retrieval-augmented recommendation is proposed, where the most relevant information is retrieved from the historical data. Specifically, UBR4CTR (Qin et al., 2020) and SIM (Ren et al., 2019) are designed to retrieve beneficial behaviors from the user’s historical extremely long behavior sequence, where UBR4CTR deploys the search engine method while SIM uses the hard search and soft search approaches. To make the search procedure end-to-end, ETA (Chen et al., 2022) is proposed by leveraging the SimHash algorithm to map the user behavior into a low-dimensional space, hence achieving learnable retrieval. Moreover, recent works further extend the retrieval-augmented recommendation from item-level retrieval to sample-level retrieval by retrieving informative samples. RIM (Qin et al., 2021) is the first to deploy this method that leverages the search engine to retrieve several relevant samples from the search pool and performs neighbor aggregation. PET (Du et al., 2022) and DERT (Zheng et al., 2023) have respectively made improvements in the interaction and retrieval mechanism of neighboring samples. PET constructs a hypergraph over the retrieved data samples and performs message propagation to improve the target data representations for final CTR prediction. DERT utilizes vector retrieval to speed up the retrieval process.
3. Preliminary
3.1. CTR Prediction
[…]
Apart from the traditional training/validation/testing splits for CTR prediction, retrieval-based methods further require a search pool $T_{pool}$, which might overlap with the training set $T_{train}$ according to different settings. The search pool $T_{pool}$ is constructed to provide useful knowledge for downstream CTR prediction towards the target sample $x_{t}$, which can be formulated as:
(1) $\hat{y}_{t}=G\left(x_{t},R\left(x_{t}\right);\mathbf{\theta}\right),$
where $R\left(x_{t}\right)$ is the retrieved knowledge of $x_{t}$. For item-level retrieval (Qin et al., 2020; Pi et al., 2020), the knowledge $R\left(x_{t}\right)$ is the retrieved $k$ user behaviors. For sample-level retrieval (Qin et al., 2021; Du et al., 2022), the knowledge $R\left(x_{t}\right)$ is the retrieved $k$ nearest data samples. As for our proposed ROK, the knowledge $R\left(x_{t}\right)$ is the learned retrieval-enhanced representation.
After obtaining the click prediction $\hat{y}_{t}$, the parameters $\theta$ are optimized by minimizing the binary cross-entropy (BCE) loss:
(2) $\mathcal{L}_{t}=y_{t}\log\hat{y}_{t}+(1-y_{t})\log(1-\hat{y}_{t}).$
3.2. Sample-level Retrieval-based Methods
For illustration purposes, we abstract the sample-level retrieval approach in the left part of Figure 2. The target sample $x_{t}$ will be considered as a query to retrieve the top-K neighboring samples $\{s_{1},\cdots,s_{K}\}$ from the search pool $T_{pool}$ through the search engine. After retrieval, an aggregation layer is used to aggregate the features and labels of the retrieved samples, resulting in a dense knowledge representation $\mathbf{r}_{t}$. For example, RIM (Qin et al., 2021) deploys an attentive aggregation layer, while PET (Du et al., 2022) constructs a hypergraph over the retrieved samples, and performs message propagation to aggregate the representations. Finally, the aggregated representation $\mathbf{r}_{t}$ is sent to the prediction module together with the target sample representation. The prediction module usually contains components for feature interaction modeling (Chen et al., 2021b; Guo et al., 2017; Song et al., 2019) or user behavior modeling (Zhou et al., 2019, 2018).
4. Methodology
4.1. Overview of ROK
Despite the superior performance, sample-level retrieval-based methods suffer from the inference inefficiency problem due to the time-consuming online retrieval process. To this end, we propose a novel Retrieval-Oriented Knowledge (ROK) framework, where a knowledge base is built to imitate the aggregated retrieval knowledge $\mathbf{r}_{t}$. As shown in Figure 2, ROK consists of two stages: (1) Retrieval-Oriented Knowledge Construction, and (2) Knowledge Utilization.
In the first stage of ROK–retrieval-oriented knowledge construction–we pre-train a sample-level retrieval-based model [^2] In this work, we employed RIM(Qin et al., 2021) which is one of the foundational works in sample-level retrieval-based methods. The pre-training process exactly follows the original work. Importantly, our proposed method extends its applicability to encompass all sample-level retrieval-based methods., followed by designing a delicate knowledge base with a retrieval-oriented embedding layer and a knowledge encoder. This is designed to imitate the retrieval & aggregation results (i.e., $\mathbf{r}_{t}$) of the pre-trained model by directly generating the final aggregated representation $\mathbf{z}_{t}^{x}$ through the knowledge encoder. In this way, when online serving, the time-consuming retrieval process could be replaced by a simpler and faster forward propagation of a neural network (i.e., knowledge base). We adopt the mean square error loss for knowledge imitation and design a contrastive regularization loss to stabilize the learning process and prevent collapse (Jing et al., 2021; Shen et al., 2022).
In the knowledge utilization stage, we propose to integrate the plug-and-play retrieval-enhanced representations into arbitrary backbone CTR models in both instance-wise and feature-wise manner. Hence, we retain the sample-level retrieval capacity, while avoiding the online overhead brought by retrieval & aggregation operations.
Next, we will first elaborate on the network structure design of the knowledge base for knowledge imitation. Then, we give the detailed training strategies for the two stages (i.e., retrieval-oriented knowledge construction and knowledge utilization).
4.2. Structure Design of Knowledge Base
For sample-level retrieval-based methods (Qin et al., 2021; Du et al., 2022), the overhead of inference inefficiency is mainly caused by the retrieval & aggregation operations, which are uncacheable and unavoidable during the online inference if a brand-new data sample comes. To this end, we propose a novel decomposition-reconstruction paradigm, where the aggregated representation $\mathbf{r}_{t}$ is first decomposed into feature-level embeddings, and then reconstructed via a neural encoder. As illustrated in  3, our proposed knowledge base comprises two key modules: (1) retrieval-oriented embedding layer $f$, and (2) knowledge encoder $g$.
Initially, we decompose the query $x_{t}$ into feature-wise embeddings via the retrieval-oriented embedding layer $f$. Then, the obtained feature-wise embeddings are fed into the knowledge encoder $g$ to reconstruct the instance-wise aggregated representation $\mathbf{z}_{t}^{x}=g(f(x_{t}))$. Specifically, a learnable embedding layer functions as the retrieval-oriented embedding layer. The architecture of knowledge encoder $g$ can be chosen arbitrarily (e.g., transformer (Vaswani et al., 2017)). For simplicity, we adopt multi-layer perceptron (MLP) in this paper. This approach reduces the time complexity of sample-level retrieval from $\mathcal{O}(N\log\left(N\right))$, where $N$ is the search pool size to $\mathcal{O}(1)$. Since we introduce the retrieval-oriented embedding layer $f$, we will cut down the embedding size of $f$ from $d$ to $d/2$ for fair comparison in space complexity. With the help of the learned knowledge base, online retrieval & aggregation operations can be avoided.
Figure 3. Structure Design of Knowledge Base.
4.3. Retrieval-Oriented Knowledge Construction
To construct the knowledge base, we introduce two modules: Retrieval Imitation and Contrastive Regularization.
4.3.1. Retrieval Imitation
After we obtain the reconstructed representation $\mathbf{z}_{t}^{x}$ from the knowledge base for query sample $x_{t}$. We use mean square error (MSE) loss to train the knowledge base, enabling it to imitate the aggregated representation $\mathbf{r}_{t}$:
(3) $\mathcal{L}_{imit}=\mbox{MSE}\left(\mathbf{z}^{x}_{t},\mathbf{r}_{t}\right).$
This knowledge imitation approach serves as a knowledge distillation process, where we extract and inject the retrieval-enhanced knowledge from the pre-trained model into the learnable knowledge base, enabling the implicit sample-level retrieval capability.
4.3.2. Contrastive Regularization
Besides, we have intricately integrated a contrastive regularization loss to ensure stability in the learning process and avert potential collapse (Jing et al., 2021; Shen et al., 2022). Notably, sample-level retrieval-based methods (Qin et al., 2021; Du et al., 2022) ingeniously produces positive samples during the retrieval process. This can be perceived as a unique data augmentation technique, offering a localized positive perspective for contrastive learning. Consequently, ROK employs the SimSiam (Chen and He, 2021) framework, utilizing a free negative samples scheme for regularization. While the traditional sample-level retrieval-based methods predominantly harness global features from the aggregated neighboring samples, our use of contrastive regularization allows for a more detailed extraction of local features from these samples.
As shown in Figure 2(a), the most correlated neighboring sample $s_{t}$ is selected from the $K$ retrieved samples $\{s_{1},\cdots,s_{K}\}$ by the sample-level retrieval-based methods. Then, the target sample $x_{t}$ and selected neighboring sample $s_{t}$ are fed into the knowledge base to obtain the reconstructed representations $\mathbf{z}^{x}_{t}$ and $\mathbf{z}^{s}_{t}$, respectively. To prevent collapse in the absence of negative samples (Zhang et al., 2022), a projector $h$ is employed to generate the projected representation $\mathbf{p}^{x}_{t}$, $\mathbf{p}^{s}_{t}$ from $\mathbf{z}^{x}_{t}$, $\mathbf{z}^{s}_{t}$.
(4) $\vspace*{-0.5\baselineskip}\mathbf{p}^{x}_{t}$ $\triangleq h\left(\mathbf{z}^{x}_{t}\right)\triangleq h\left(g\left(f\left(x_{t}\right)\right)\right),$
(5) $\mathbf{p}^{s}_{t}$ $\triangleq h\left(\mathbf{z}^{s}_{t}\right)\triangleq h\left(g\left(f\left(s_{t}\right)\right)\right).\vspace*{-0.5\baselineskip}$
The cosine similarity between projected representation $\mathbf{p}^{x}_{t}$ and reconstructed representations $\mathbf{z}^{s}_{t}$ is defined as:
(6) $\mathcal{D}\left(\mathbf{p}^{x}_{t},\mathbf{z}^{s}_{t}\right)=\frac{\mathbf{p}^{x}_{t}}{\left\|\mathbf{p}^{x}_{t}\right\|_{2}}\cdot\frac{\mathbf{z}^{s}_{t}}{\left\|\mathbf{z}^{s}_{t}\right\|_{2}},\vspace*{-0.5\baselineskip}$
where $\left\|\cdot\right\|_{2}$ is the $l_{2}$ norm. Besides, the gradient of $\mathbf{z}^{s}_{t}$ is stopped when computing the loss (Zhang et al., 2022; Chen and He, 2021). To avoid spatial bias, motivated by the Jensen–Shannon (JS) divergence, a symmetric loss is adopted and the final contrastive regularization loss is:
(7) $\mathcal{L}_{contra}=-\left(\frac{1}{2}\mathcal{D}\left(\mathbf{p}^{x}_{t},\mathbf{z}^{s}_{t}\right)+\frac{1}{2}\mathcal{D}\left(\mathbf{p}^{s}_{t},\mathbf{z}^{x}_{t}\right)\right).\vspace*{-0.5\baselineskip}$
Hence, The overall objective for ROK in the retrieval-oriented knowledge construction stage is:
(8) $\mathcal{L}=(1-\alpha)\cdot\mathcal{L}_{imit}+\alpha\cdot\mathcal{L}_{contra},\vspace*{-0.5\baselineskip}$
where $\alpha$ adjusts the ratio of the two loss terms. After the retrieval-oriented knowledge construction stage, the knowledge base will be retained for the follow-up knowledge utilization stage.
4.4. Knowledge Utilization
In the knowledge utilization stage, the plug-and-play retrieval-enhanced representations from the learned knowledge base can be integrated into various backbone CTR models. When integrated with backbone models, all parameters of ROK are frozen—a strategy that outperforms others, as detailed in Section 5.5.1. As illustrated in Figure 2(b), we present two approaches for knowledge utilization, which are mutually compatible to boost the final CTR prediction:
  • Feature-wise retrieval-enhanced method leverages the retrieval-oriented embedding layer $f$ to obtain the retrieval-enhanced feature-wise embeddings for the target sample. Specifically, this method integrates the retrieval-oriented embedding layer $f$ with the original learnable feature embedding of the backbone model to form the final embedding layer. 333 The embedding layer accounts for a large proportion of the total number of parameters in CTR models. For a fair comparison in model parameters, the embedding size of both the retrieval-oriented embedding layer and the original learnable feature embedding of the backbone model is set to $d/2$ compared with other models. .
  • Instance-wise retrieval-enhanced method employs the knowledge base to generate a retrieval-enhanced instance-level aggregated representation $\mathbf{z}_{t}^{x}$ for each sample. Typically, various backbone models produce their unique representations—for example, RIM’s aggregated representations and DIEN’s interest states. These unique representations are then concatenated with the feature embeddings and the combined results are fed into the interaction and MLP (Multi-Layer Perceptron) layers.
    $\left\{\begin{array}[]{ll}\mathbf{z}_{t}^{x}&=g(f(x_{t}))\\ \hat{y}_{t}&=\text{backbone}(\phi(x_{t})\oplus\mathbf{z}_{t}^{x})\end{array},\right.$
    where $\phi$ is utilized to extract the unique representations of the backbone model.
5. Experiments
In this section, we present the experimental settings and corresponding results in detail. The experiments are conducted on three large-scale datasets, including Tmall, Taobao, and Alipay. To gain more insights into ROK, we endeavor to address the following research questions (RQs) in this section.
  • RQ1: How does ROK’s performance compare to that of traditional CTR models and retrieval-based methods?
  • RQ2: How compatible is ROK with other backbone models?
  • RQ3: How does the knowledge captured by ROK contribute to improving performance?
  • RQ4: How does the chosen update strategy and positive sample selecting strategy affect ROK’s performance?
  • RQ5: How do the specified hyperparameter influence ROK’s performance?
5.1. Experimental Settings
5.1.1. Datasets.
The evaluations were conducted on three widely-recognized public datasets: Tmall, Taobao, and Alipay. Comprehensive statistics for these datasets are presented in Table 1. Specifically, we tallied the counts of users, items, samples, fields, and categories for each dataset. It’s noteworthy to mention that the number of features refers to the count of unique feature values.
Table 1. The dataset statistics.
Dataset Users # Items # Samples # Fields # Features #
Tmall 424,170 1,090,390 54,925,331 9 1,529,676
Taobao 987,994 4,162,024 100,150,807 4 5,159,462
Alipay 498,308 2,200,291 35,179,371 6 3,327,205
Following RIM (Qin et al., 2021), we organized the data such that the oldest instances constituted the search pool, the most recent instances comprised the test set, and those in between were assigned to the training set. The retrieval-based methods (Qin et al., 2021, 2020) retrieve neighboring samples from the search pool. For user behavior modeling methods (Zhou et al., 2018, 2019) involved (e.g., DIEN), the sequential features (e.g., user behavior) are also generated from the search pool.
5.1.2. Evaluation Metrics.
To measure the performance, we utilize the commonly adopted metrics AUC and log-loss, which reflect pairwise ranking performance and point-wise likelihood, respectively. A significance test contrasting the two top-performing methods for each metric is undertaken, with results of significance denoted by an asterisk (*).
We follow (Zhou et al., 2018) to introduce Rel. Impr. metric to quantify the relative improvement of models, which is defined as below:
[…]
where the base model is the backbone model of ROK in each dataset.
5.1.3. Compared Baselines.
For baselines, we compare ROK with a mid-tier performance backbone model against traditional models and retrieval-based models. For traditional CTR models, GBDT (Chen et al., 1996) is a frequently utilized tree-based model, and DeepFM (Guo et al., 2017) is a popular deep-learning model that focuses on feature interactions. HPMN (Ren et al., 2019) and MIMN (Pi et al., 2019) utilize memory network architectures, while DIN (Zhou et al., 2018) and DIEN (Zhou et al., 2019) are attention-based CTR models designed to pinpoint user interests. Additionally, FATE (Wu et al., 2021) is a model for learning representations of tabular data which facilitates interactions among samples within a minibatch. In item-level retrieval-based models, SIM (Pi et al., 2020) and UBR (Qin et al., 2020) extract pertinent user behaviors from a comprehensive set of user-generated data. As for the sample-level retrieval-based model, RIM (Qin et al., 2021) retrieves relevant data instances based on raw input features, and DERT (Zheng et al., 2023) retrieves dense representations. To ensure a fair comparison, we adopted the same hyper-parameter settings for the baselines and the ROK backbone during the knowledge utilization phase as those used in prior work (Qin et al., 2021).
Regarding backbone models, we intentionally select average-performing common models like DIN and DIEN rather than top-tier ones like UBR, which will further improve the performance of ROK. By default, we choose DIEN (Zhou et al., 2019) for Tmall and Taobao datasets and DIN (Zhou et al., 2018) for the Alipay dataset. The substantial improvement observed in 5.2 when these backbone models are combined with ROK underscores the efficacy of ROK.
5.2. Overall Performance Comparison: RQ1
In this section, we compare the overall experimental results in Table 2, from which we have the following observations. First, ROK remarkably boosts the backbone model with AUC improved by 10.08%, 34.96%, and 16.85% on the three datasets respectively, demonstrating the superior performance. Second, retrieval-based methods (Qin et al., 2020, 2021) significantly outperform user behavior modeling methods (Zhou et al., 2018, 2019), benefiting from their superior knowledge retrieval ability. Moreover, sample-level retrieval-based methods (Qin et al., 2021) perform better than item-level retrieval-based methods (Pi et al., 2020; Qin et al., 2020) due to more comprehensive retrieval. Finally, ROK markedly outperforms the item-level retrieval-based methods such as SIM (Pi et al., 2020) and UBR (Qin et al., 2020), and achieves performance surpassing that of the sample-level retrieval-based methods like RIM (Qin et al., 2021). Although ROK is only based on knowledge distillation and contrastive learning, it not only surpasses the teacher model but also comes very close to recent results in terms of performance. Notably, in our methodology, we deliberately opted for models of average performance, such as DIN and DIEN, over elite models like UBR (Qin et al., 2020), with the anticipation that the latter would enhance the efficacy of ROK. Concurrently, it is pertinent to note that the teacher model employed within ROK is RIM, which DERT outperforms. Nonetheless, the adoption of a superior teacher model is poised to markedly augment the performance of ROK. Despite this, ROK has successfully enabled user behavior modeling methods to outperform both item-level retrieval-based methods and sample-level retrieval-based methods. This indicates that the neural knowledge model successfully serves as a compact surrogate for the retrieval pool. More importantly, the sample-level retrieval-based methods, due to their prolonged inference time, were once deemed entirely infeasible for industrial application, which will further be elaborated in Section 6.2. However, with the introduction of ROK, this barrier has been overcome. ROK not only mitigates the challenges but revolutionizes the deployment prospects, rendering sample-level retrieval-based methods feasible industry applications.
Table 2. Performance comparison of CTR prediction task with various baselines. For each dataset, RIM is selected as the teacher model, and an average-performing backbone model is selected for ROK: DIEN for both the Tmall and Taobao datasets and DIN for the Alipay dataset. indicates recent results. Evaluation metrics include AUC and log-loss (LL). Among the models excluding recent work, the best results are highlighted in bold, while the runner-ups are underlined. The gray-colored bold text indicates recent results outperform all other models. “Rel. Impr.” signifies the model’s relative AUC improvement over the chosen backbone model, with a statistical significance level of $p<0.01$.
Category Model Tmall Taobao Alipay
AUC LL Rel. Impr AUC LL Rel. Impr AUC LL Rel. Impr
Traditional Models GBDT 0.8319 0.5103 -13.55% 0.6134 0.6797 -54.75% 0.6747 0.9062 -34.00%
DeepFM 0.8581 0.4695 -6.72% 0.671 0.6497 -31.76% 0.6971 0.6271 -25.54%
FATE 0.8553 0.4737 -7.45% 0.6762 0.6497 -29.69% 0.7356 0.6199 -10.99%
HPMN 0.8526 0.4976 -8.15% 0.7599 0.5911 3.71% 0.7681 0.5976 1.28%
MIMN 0.8457 0.5008 -9.95% 0.7533 0.6002 1.08% 0.7667 0.5998 0.76%
DIN 0.8796 0.4292 -1.12% 0.7433 0.6086 -2.91% 0.7647 0.6044 0.00%
DIEN 0.8839 0.4272 0.00% 0.7506 0.6082 0.00% 0.7485 0.6019 -6.12%
Retrieval-based Models SIM (Item-level) 0.8857 0.4520 0.47% 0.7825 0.5795 12.73% 0.7600 0.6089 -1.78%
UBR (Item-level) 0.8975 0.4368 3.54% 0.8169 0.5432 26.46% 0.7952 0.5747 11.52%
RIM (Teacher model) 0.9151 0.3697 8.13% 0.8567 0.4546 42.34% 0.8005 0.5736 13.52%
[…]
Our Model
ROK 0.9226 0.3546 10.08% 0.8382 0.5098 34.96% 0.8093 0.5304 16.85%
5.3. Compatibility Analysis: RQ2
In this section, we commence by contrasting the compatibility of our novel approach, ROK, with conventional CTR models, with the findings delineated in Table 3. Upon examining the data, two observations arise: (i) The application of ROK markedly amplifies the performance of the backbone models. To illustrate, by integrating ROK with DeepFM, DIN, and DIEN, there’s a notable enhancement in performance metrics, with an average AUC increase of 4.30%, 8.01%, and 5.59%, coupled with an average decrease in log-loss by 11.80%, 8.51%, and 8.22% across three datasets, respectively. This highlights the significant benefits of incorporating the retrieval-oriented knowledge from ROK into the backbone models. (ii) Exhibiting model-agnosticism, ROK seamlessly integrates with a diverse array of backbone CTR models, irrespective of whether they focus on feature interaction or behavior modeling. This underscores that the knowledge garnered by ROK is beneficial across diverse models, highlighting the compatibility of ROK. The feature of model-agnosticism is crucial since, regardless of the optimal model for a given scenario, ROK consistently boosts its capabilities. Thus, instead of being bound to one model, we can choose the ideal one for each situation and leverage ROK to optimize its performance.
Table 3. Compatibility analysis of applying ROK to different backbone models. “Rel. Impr.” means relative AUC and log-loss improvements of ROK against the backbone. Improvements are statistically significant with $p<0.01$.
Model Tmall Taobao Alipay
AUC Rel. Impr. LL Rel. Impr. AUC Rel. Impr. LL Rel. Impr. AUC Rel. Impr. LL Rel. Impr.
DeepFM 0.8585 4.21% 0.4803 9.31% 0.6710 3.52% 0.6497 2.51% 0.6971 5.59% 0.6271 4.80%
DeepFM+ROK 0.8946 - 0.4356 - 0.6946 - 0.6334 - 0.7361 - 0.5970 -
DIN 0.8796 4.26% 0.4292 5.59% 0.7433 8.85% 0.6086 6.82% 0.7647 5.83% 0.6044 12.24%
DIN+ROK 0.9171 - 0.4052 - 0.8091 - 0.5671 - 0.8093 - 0.5304 -
DIEN 0.8839 4.38% 0.4272 20.50% 0.7506 11.67% 0.6082 16.18% 0.7485 5.33% 0.6019 7.61%
DIEN+ROK 0.9226 - 0.3546 - 0.8382 - 0.5098 - 0.7884 - 0.5561 -
5.4. Comparison of the learned knowledge of ROK and RIM: RQ3
In this section, we delve into the comparative analysis of the knowledge quality acquired by ROK and RIM. For a detailed assessment, we formulated two simplest variants (LR and MLP) for each model to fully demonstrate the effect of knowledge:
  • For ROK, we designed the ROK(LR) and ROK(MLP) variants. Both of these take as input the concatenation of the original feature embeddings $\mathbf{x}_{t}$, with ROK’s knowledge component $z_{t}^{x}$.
  • Correspondingly, for RIM, we introduced the RIM(LR) and RIM(MLP) variants. Their input comprises the concatenation of the original feature embeddings, $\mathbf{x}_{t}$, with RIM’s aggregated features and labels, $\mathbf{r}_{t}$.
The detailed performance comparison of these variants is tabulated in Table 4 and we have two observations. First, significantly, the results indicate that the ROK variants surpass the RIM variants in performance across both datasets. This distinction in outcomes emphasizes the superior quality of knowledge learned by ROK, especially as these models apply this intrinsic knowledge directly for predictions. It’s compelling to note that while ROK leverages the retrieval imitation strategy (detailed in Section 4.3.1) to imitate the aggregated features and labels of RIM, it extracts richer and more profound knowledge from the data. We are inclined to believe that the contrastive regularization methodology is pivotal in driving this enhanced performance for ROK. Second, ROK(MLP) in Tmall even outperforms DeepFM and some user behavior modeling methods, which shows another way to improve the pre-ranking model effect.
To delve deeper into the inherent knowledge of ROK and explain the superiority of ROK over RIM, we employed t-SNE visualization (Van der Maaten and Hinton, 2008). Figure 4 and Figure 7 depict the knowledge distribution patterns of both ROK and RIM across the Alipay and Tmall datasets. For this visualization, we randomly selected 10,000 samples from both the training and testing sets for each model. We have two observations. First, a discerning observation from Figure 4 reveals that ROK’s knowledge distribution exhibits a more pronounced clustering effect across all datasets. This enhanced clustering can be attributed to contrastive regularization. For comparison, RIM’s retrieved neighboring samples may include some incidental noise that correlates with the quantity of these samples. Besides, the aggregation mechanism of RIM only captures the global features of the neighboring samples. However, the contrastive regularization in ROK emphasizes the local features of neighboring samples and mitigates the noise of neighboring samples by selectively considering the most relevant sample, as shown in section 4.3.2. Specifically, representations of analogous samples are pulled closer, while those of dissimilar samples are pushed apart, leading to a more distinct and meaningful clustering. Secondly, a significant concern when using neural networks as a knowledge base in our case relates to their generalization abilities. This is because the training process is guided by the retrieved examples of sample-level retrieval-based methods, while the testing phase requires the model to apply its knowledge independently. Nonetheless, the visualization of ROK’s representations reveals a high degree of similarity between the training and test datasets, showing the good generalization of the decomposition-reconstruction paradigm.
Table 4. Comparison of model variants. “Rel. Impr.” means relative AUC improvement of ROK’s variants against baselines and RIM’s variants. Improvements are statistically significant with $p<0.01$.
Model Tmall Alipay
AUC Rel. Impr. AUC Rel. Impr.
LR 0.8213 4.37% 0.6298 8.96%
RIM(LR) 0.8379 2.30% 0.6613 3.77%
ROK(LR) 0.8572 - 0.6862 -
MLP 0.8393 2.98% 0.6344 9.13%
RIM(MLP) 0.8426 2.58% 0.6773 2.21%
ROK(MLP) 0.8643 - 0.6923 -
Refer to captionRefer to caption
((a)) Alipay (Train)
Refer to captionRefer to caption
((b)) Alipay (Test)
Figure 4. The t-SNE visualization of knowledge from RIM (on the left in each subfigure) and ROK (on the right in each subfigure) for Alipay.
5.5. Ablation & Hyper-parameter Study
5.5.1. Update Strategy: RQ4.
In this section, we undertake comprehensive ablation experiments for ROK. We start by examining the parameter update strategy employed for the knowledge base during the knowledge utilization phase. The strategies under consideration include: fixing the knowledge base (Fix), updating only the upper knowledge encoder $g$ (Upd $g$), and updating both the knowledge encoder and the retrieval-oriented embedding layer $f$ (Upd $f+g$). Furthermore, we consider the positive sample selection strategy mentioned in Section 4.3.2. Specifically, we select the most related neighboring sample by default. Now, we compare the performance of randomly selecting a sample from the retrieved samples (random). Results obtained on the Tmall and Alipay datasets are depicted in Figure 5. We have two observations: First, we observe that randomly selecting a sample as the positive sample diminishes performance. This suggests that even within the narrowly retrieved set of samples, noise is still present. Second, a key observation from these results is that the strategy of fixing the knowledge base (Fix) yields the best performance. This can be attributed to the primary goal of ROK: to derive superior representations. However, updating with the backbone models may adversely impact the acquired knowledge. This characteristic is highly desirable and aligns well with our expectations, further enhancing the practicality of ROK. This indicates that once ROK has been fully trained, it can function as an independent module. This eliminates the need for subsequent modifications, reinforcing its role as an authentic knowledge base.
((a)) Tmall
((b)) Alipay
Figure 5. Comparison of AUC scores for different update strategies.
5.5.2. Hyperparameter Study: RQ5.
In this section, we delve into the study of the hyperparameter $\alpha$ pertaining to the balance between the contrastive regularization loss and retrieval imitation loss, represented as $\alpha$. As illustrated in Figure 6, for both datasets, when $\alpha$ assumes values of either 00 or $1$—representing the extreme scenarios—the performance of ROK deteriorates. This observation underscores the significance of both retrieval imitation loss and contrastive regularization loss to the model’s efficacy.
((a)) Tmall
((b)) Alipay
Figure 6. Comparison of AUC scores across varying $\alpha$ values.
6. Industry Application
In this section, we will demonstrate the superior performance of ROK in training and inference efficiency.
6.1. Training Efficiency
In evaluating model efficiency, it’s vital to look at both inference speed and training time. Training ROK involves pre-training a sample-level retrieval-based model, creating retrieval-oriented knowledge, and applying this knowledge. Despite these steps, the extra training time is reasonable compared to a standalone model. The training duration is detailed in Table 5. Experiments were run on an AMD EPYC 7T83 with an RTX 4090, using hyperparameters—learning rate, batch size, weight decay—optimized via grid search as per the RIM study. We also implemented early stopping after 3 steps to enhance efficiency. Two observations stand out. Firstly, incorporating ROK into the backbone model significantly reduces training time, thanks to fewer parameters needing updates, partly because ROK’s parameters are frozen and the embedding size is halved. This leads to faster convergence. Second, the total time of Phase 1 and Phase 2 is close to the training duration of the DIEN without ROK. The relatively time-consuming part of the whole training process would be the RIM’s retrieval mechanism. However, this is mitigated by using large bulk sizes and caching, which decreases retrieval times. For further efficiency, extending the knowledge base update frequency beyond the backbone model’s could be beneficial.
Table 5. Training duration comparison for two phases, total time and backbone model on Tmall, Taobao, and Alipay, measured in minutes. Phase 1 includes data retrieval for training and testing tests, pre-training RIM, and Retrieval-Oriented Knowledge Construction. Phase 2 includes Knowledge Utilization, with DIEN as the backbone model. Additionally, DIEN training times without ROK integration are provided. For Phase 1, data retrieval times are specified separately.
Phase Tmall Taobao Alipay
Phase 1 (Retrieval) 156 (40) 116 (30) 24 (4)
Phase 2 201 163 68
Total 357 279 92
DIEN without ROK 327 330 79
6.2. Inference Efficiency
In the real world of online services, from advertising platforms to recommendation systems, the significance of inference efficiency cannot be ignored. The introduction of the knowledge base significantly amplifies the efficiency in sample-level retrieval, trimming down the time complexity from $\mathcal{O}(N\log\left(N\right))$, where $N$ is the search pool size to a constant time $\mathcal{O}(1)$ through bypassing conventional retrieval & aggregation mechanisms.
Within this context, we thoroughly assess the inference efficiency of various backbone models—DeepFM (Guo et al., 2017), DIEN (Zhou et al., 2019), and DIN (Zhou et al., 2018)—when seamlessly integrated with ROK and juxtaposed against retrieval-based methods such as UBR (Qin et al., 2020), RIM (Qin et al., 2021), and DERT (Zheng et al., 2023). The comparisons drawn from our analyses, as detailed in Table 6, offer insights. Specifically, due to the $\mathcal{O}(N\log\left(N\right))$ complexity of the online retrieval process, the inference time of retrieval-based methods is significantly longer than other baseline models. For instance, RIM’s inference time is an overwhelming $1\sim 2$ orders of magnitude longer than DeepFM, DIEN, and DIN. This significant delay, despite any performance enhancements, fails to align with the real-time response expectations of RSs. These systems necessitate a swift response, making such latencies of retrieval-based methods industrially infeasible. Moreover, deploying retrieval-based methods for real-world recommendation systems remains complex and costly due to challenges such as encoding discretized features to dense representations, high resource consumption, and divergent inference processes. In contrast, our innovative ROK distinguishes itself by delivering notable improvements over the backbone models while maintaining negligible additional inference time, making sample-level retrieval-based methods feasible in online service deployment. As shown in Table 6, ROK consistently improves the AUC of backbone models across all datasets while only slightly increasing the inference speed by 0.12ms to 0.25ms per sample. Compared to retrieval-based methods, ROK achieves comparable or even better performance with significantly faster inference speed. For example, on the Tmall dataset, DIEN+ROK surpasses the performance of DERT (0.9226 vs. 0.9200) while being more than 3 times faster in inference (5.51ms vs. 17.78ms per sample).
ROK’s design allows it to be retrieval-free and compatible with existing recommendation systems. It only needs to perform inference once, avoiding additional data transfer overhead and making it highly efficient. This represents a blend of predictive accuracy and rapid inference, making ROK a noteworthy solution for modern online services.
Table 6. Comparison of AUC and inference speed (ms per sample) across models on Tmall, Taobao, and Alipay.
Model Tmall Taobao Alipay
AUC Inference Speed AUC Inference Speed AUC Inference Speed
DeepFM 0.8585 1.34 0.6710 1.36 0.6971 1.19
DeepFM+ROK 0.8946 1.46 0.6946 1.40 0.7361 1.41
DIEN 0.8839 5.44 0.7506 4.89 0.7485 3.69
DIEN+ROK 0.9226 5.51 0.8382 5.04 0.7884 3.73
DIN 0.8796 1.28 0.7433 1.32 0.7647 1.43
DIN+ROK 0.9171 1.46 0.8091 1.36 0.8093 1.56
UBR (Retrieval) 0.8975 20.71 (17.67) 0.8169 56.45 (53.32) 0.7952 30.32 (27.31)
RIM (Retrieval) 0.9151 174.81 (173.27) 0.8567 206.22 (204.78) 0.8005 113.95 (112.37)
DERT (Retrieval) 0.9200 17.78 (16.19) 0.8647 19.53 (16.97) 0.8087 17.95 (16.31)
7. Conclusion
This paper introduces the ROK framework, addressing the inefficiency challenges inherent in sample-level retrieval-based methods during inference. By developing a neural network-based knowledge base, ROK imitates and preserves the essential retrieval representations through a novel decomposition-reconstruction approach, further optimized via knowledge distillation and contrastive learning. This allows for the integration of retrieval-oriented knowledge with various CTR models, either at an instance or feature level. Through rigorous testing on three public datasets, ROK not only matches the performance of existing retrieval-based CTR models but also demonstrates superior inference efficiency. These findings highlight ROK’s potential to significantly improve the efficiency and effectiveness of CTR models.
Appendix A Appendix
A.1. T-sne visualization
Figure 7 depict the knowledge distribution patterns of both ROK and RIM in Tmall dataset.
Refer to captionRefer to caption
((a)) Tmall (Train)
Refer to captionRefer to caption
((b)) Tmall (Test)
Figure 7. The t-SNE visualization of knowledge from RIM (on the left in each subfigure) and ROK (on the right in each subfigure) for Tmall.
A.2. Ablation & Hyper-parameter Study on Taobao
In this section, we present the ablation and hyper-parameter study conducted on the Taobao dataset. The results are depicted in Figure 8, Figure 9.
Figure 8. Comparison of AUC scores for different update strategies on Taobao.
Figure 9. Comparison of AUC scores across varying $\alpha$ values on Taobao. At $\alpha=0$, only Retrieval Imitation loss is utilized, whereas at $\alpha=1$, only the Contrastive Regularization loss is applied.

### v2

@@
 Abstract.
-Click-through rate (CTR) prediction plays an important role in personalized recommendations.
-Recently, sample-level retrieval-based models (e.g., RIM) have achieved remarkable performance by retrieving and aggregating relevant samples.
-However, their inefficiency at the inference stage makes them impractical for industrial applications.
-To overcome this issue, this paper proposes a universal plug-and-play retrieval-oriented knowledge (ROK) framework.
-Specifically, a knowledge base, consisting of a retrieval-oriented embedding layer and a knowledge encoder, is designed to preserve and imitate the retrieved & aggregated representations in a decomposition-reconstruction paradigm.
-Knowledge distillation and contrastive learning methods are utilized to optimize the knowledge base, and the learned retrieval-enhanced representations can be integrated with arbitrary CTR models in both instance-wise and feature-wise manners.
-Extensive experiments on three large-scale datasets show that ROK achieves competitive performance with the retrieval-based CTR models while reserving superior inference efficiency and model compatibility.
-copyright: acmcopyrightjournalyear: 2023doi: XXXXXXX.XXXXXXXconference: xx; Preprint; price: xisbn: xxx
+Click-through rate (CTR) prediction is crucial for personalized online services.
+Sample-level retrieval-based models, such as RIM, have demonstrated remarkable performance.
+However, they face challenges including inference inefficiency and high resource consumption due to the retrieval process, which hinder their practical application in industrial settings.
+To address this, we propose a universal plug-and-play retrieval-oriented knowledge (ROK) framework that bypasses the real retrieval process.
+The framework features a knowledge base that preserves and imitates the retrieved & aggregated representations using a decomposition-reconstruction paradigm.
+Knowledge distillation and contrastive learning optimize the knowledge base, enabling the integration of retrieval-enhanced representations with various CTR models.
+Experiments on three large-scale datasets demonstrate ROK’s exceptional compatibility and performance, with the neural knowledge base serving as an effective surrogate for the retrieval pool.
+ROK surpasses the teacher model while maintaining superior inference efficiency and demonstrates the feasibility of distilling knowledge from non-parametric methods using a parametric approach.
+These results highlight ROK’s strong potential for real-world applications and its ability to transform retrieval-based methods into practical solutions.
+Our implementation code is available to support reproducibility [^1] https://github.com/HSLiu-Initial/ROK.git.
+Information Retrieval; Recommender Systems; Knowledge Distillation; Contrastive Learning
+journalyear: 2024copyright: acmlicensedconference: Proceedings of the 33rd ACM International Conference on Information and Knowledge Management; October 21–25, 2024; Boise, ID, USA.booktitle: Proceedings of the 33rd ACM International Conference on Information and Knowledge Management (CIKM ’24), October 21–25, 2024, Boise, ID, USAisbn: 979-8-4007-0436-9/24/10doi: 10.1145/3627673.3679842ccs: Information systems Recommender systemsccs: Information systems Nearest-neighbor searchccs: Information systems Users and interactive retrieval
 1. Introduction
-Click-through rate (CTR) prediction is a key component of many personalized online services, such as recommender systems (Xi et al., 2023) and web search (Lin et al., 2021; Fu et al., 2023; Dai et al., 2021).
-It seeks to estimate the probability of a user’s click given a particular context (Zhang et al., 2021).
-The CTR models could be mainly classified into two categories.
+Click-through rate (CTR) prediction is a key component of many personalized online services, such as recommender systems (Xi et al., 2023; Wang et al., 2023) and web search (Lin et al., 2021; Fu et al., 2023; Dai et al., 2021; Liu et al., 2024).
+It aims to estimate the probability of a user’s click given a particular context (Zhang et al., 2021).
+The CTR models can be mainly classified into two categories.
 The first category is the feature interaction-based methods.
@@
 The second category is the user behavior modeling methods that leverage different architectures (e.g., RNN (Hidasi and Karatzoglou, 2018; Hidasi et al., 2015), CNN (Tang and Wang, 2018), attention (Zhou et al., 2019, 2018), memory bank (Pi et al., 2019; Ren et al., 2019)) to extract informative knowledge from user behavior sequences for final CTR prediction.
-To further enhance the performance, UBR4CTR (Qin et al., 2020) and SIM (Pi et al., 2020) retrieve useful behaviors from the user’s behavior history (i.e., clicked items), reducing the potential noise in user behavior sequences.
-Subsequent studies (Li et al., 2022b) have improved these methods’ efficiency using hashing functions (Chen et al., 2021a) and parallel retrieval execution (Cao et al., 2022) during inference.
-Recent studies (Qin et al., 2021; Du et al., 2022) further expand these retrieval-based methods from item-level retrieval to sample-level retrieval, applicable to general CTR prediction settings.
+Due to the problem of distribution shift and the daily generation of vast amounts of interaction data in industrial applications, using all the data for training becomes impractical.
+This results in a significant waste of historical data, as traditional methods cannot effectively utilize it.
+To address this issue, retrieval-based methods have been proposed.
+UBR4CTR (Qin et al., 2020) and SIM (Pi et al., 2020) retrieve useful behaviors from the user’s behavior history (i.e., clicked items), reducing the potential noise in user behavior sequences.
+Subsequent studies (Li et al., 2022b) improve these methods’ efficiency using hashing functions (Chen et al., 2021a) and parallel retrieval execution (Cao et al., 2022) during inference.
+Recent studies (Qin et al., 2021; Du et al., 2022) further expand these retrieval-based methods from item-level retrieval to sample-level retrieval, making them applicable to general CTR prediction settings.
 Instead of retrieving similar items from the user history, RIM (Qin et al., 2021) adopts the idea of $k$ nearest neighbor ($k$NN) and designs a sample-centric retrieval method, which aggregates the relevant data samples retrieved from the search pool (e.g., the whole training dataset).
-PET (Du et al., 2022) constructs a hypergraph over the retrieved data instances and performs message propagation to get a better representation of the target data for final CTR prediction.
-Although sample-level retrieval-based methods bring impressive performance enhancement, they have to perform instance-wise comparisons between the target data sample and each candidate sample in the search pool (usually million or even billion level).
-This leads to extreme inefficiency problems during inference, making it impractical for industrial applications, as illustrated in the left part of Figure 1.
-Recent work, such as DERT (Zheng et al., 2023), has improved the retrieval mechanisms by using vector retrieval techniques, enhancing efficiency.
-However, DERT’s dependency on the RIM Encoder and its $O(\log N)$ retrieval time complexity, along with the additional cost of encoding numerous samples in the retrieval pool, limits its scalability and effectiveness in large-scale applications.
-This work introduces the Retrieval-Oriented Knowledge (ROK) framework to tackle the inference inefficiency problem of sample-level retrieval-based methods, with two stages: Retrieval-Oriented Knowledge Construction and Knowledge Utilization.
-Initially, we pre-train a sample-level retrieval-based model, such as RIM(Qin et al., 2021).
-To resolve the inefficient inference issue, ROK tries to learn a neural network-based Knowledge Base to imitate the aggregated representations from the pre-trained retrieval-based methods via a Retrieval Imitation module, whose simple diagram is shown in Figure 1.
+Although sample-level retrieval-based methods brings impressive performance enhancement, they have to perform instance-wise comparisons between the target data sample and candidate samples in the search pool (usually at the million or even billion level), leading to extreme inefficiency problems during inference, as illustrated in the left part of Figure 1.
+While follow-up studies (Zheng et al., 2023; Li et al., 2024) have attempted to enhance retrieval mechanisms through vector retrieval techniques, they have instead created further challenges and failed to meet the necessary inference speed for deployment.
+Moreover, sample-level retrieval-based methods faces several other significant challenges.
+These include high resource consumption due to the linear increase in retrievers with request volume, as well as a divergent inference process that requires an inference-retrieval-inference cycle.
+This necessitates significant optimizations, which makes it challenging for industrial applications.
+This work introduces the Retrieval-Oriented Knowledge (ROK) framework to address the problems of sample-level retrieval-based methods by bypassing the real retrieval process.
+The framework operates in two stages: Retrieval-Oriented Knowledge Construction and Knowledge Utilization.
+In the first stage, we pre-train a sample-level retrieval-based method, such as RIM (Qin et al., 2021).
+To bypass the real retrieval process, ROK learns a neural network-based Knowledge Base to store the retrieval-oriented knowledge from this pre-trained method.
+Specifically, ROK imitates the aggregated representations from the pre-trained retrieval-based method via a Retrieval Imitation module, as illustrated in Figure 1.
 The Knowledge Base, built on a novel decomposition-reconstruction paradigm, where a Retrieval-Oriented Embedding Layer captures the feature-wise embedding and a Knowledge Encoder reconstructs the instance-wise aggregated representations (i.e., retrieval-enhanced representation).
-In this way, instead of the time-consuming retrieval, efficient inference of neural networks could be adopted to get the approximated aggregated representations.
+In this way, instead of the time-consuming retrieval, efficient inference of neural networks can be adopted to get the approximated aggregated representations.
 Additionally, we introduce a Contrastive Regularization module to ensure learning stability and prevent model collapse.
@@
 The main contributions of this paper are as follows:
-  • In this work, we introduce ROK, a pioneering framework that transforms sample-level retrieval-based methods —previously deemed infeasible in the industry into a practical solution.
-Our approach innovatively leverages a neural network-based Knowledge Base to efficiently store the retrieval-enhance representations and hence obviates the need for time-intensive retrieval & aggregation during inference.
-  • We use knowledge distillation and contrastive learning to optimizie the knowledge base, allowing the integration of retrieval-enhanced representations with various CTR models, both at the instance and feature levels.
-  • Extensive experiments conducted across three large-scale datasets have shown that ROK not only achieves competitive performance when compared to existing retrieval-based CTR methods but also reserves superior inference efficiency.
-Moreover, ROK enhances the performance of various CTR methods due to its exceptional compatibility.
-This success underscores the effectiveness of the neural knowledge model as a compact surrogate for the retrieval pool.
+  • ROK is a novel framework that leverages knowledge distillation to efficiently utilize large amounts of historical data and retrieval-oriented knowledge for training and lightweight deployment.
+To the best of our knowledge, ROK is the first such framework.
+It transforms sample-level retrieval-based methods into a practical solution by using a neural network-based Knowledge Base to store retrieval-enhanced representations, eliminating the need for time-intensive retrieval during inference.
+This demonstrates the feasibility of distilling knowledge from non-parametric models and learning their inherent knowledge using a parametric approach.
+  • ROK optimizes the Knowledge Base using a combination of knowledge distillation and contrastive learning.
+This approach allows for the seamless integration of retrieval-enhanced representations with various CTR models at both the instance and feature levels, enhancing their performance and compatibility.
+  • Extensive experiments across three large-scale datasets demonstrate ROK’s exceptional compatibility and performance.
+The results show that ROK significantly enhances the performance of various CTR methods, indicating that the neural Knowledge Base serves as an effective and compact surrogate for the search pool.
+Although the knowledge distillation foundation initially suggests that ROK may only approach the teacher model’s performance, the introduction of contrastive regularization enables ROK to surpass the sample-level retrieval-based teacher model.
+This highlights ROK’s strong potential for real-world applications.
 Figure 1. Comparison between traditional sample-level retrieval framework and ROK.
@@
 A tailored knowledge base is then designed to imitate the aggregation results, with Contrastive Regularization ensuring stabilization during the training phase.
-This design allows the time-consuming retrieval process to be substituted with the swift forward propagation of a neural network.
-In the subsequent Knowledge Utilization stage, the framework introduces the retrieval-enhanced representation into backbone CTR models in both an instance-wise and feature-wise approach, ensuring the maintenance of sample-level retrieval abilities while circumventing the complexities of retrieval and aggregation operations.
+This design allows the time-consuming retrieval & aggregration process to be substituted with the swift forward propagation of a neural network.
+In the subsequent Knowledge Utilization stage, the framework introduces the retrieval-enhanced representation into backbone CTR models in both an instance-wise and feature-wise approach, ensuring the maintenance of sample-level retrieval abilities while circumventing the complexities of retrieval & aggregation operations.
 2. Related Work
 2.1. CTR Prediction
-For the click-through rate (CTR) prediction, models can essentially be divided into two main categories: those focusing on feature interaction and those centering on user behavior modeling.
+For click-through rate (CTR) prediction, models can essentially be divided into two main categories: those focusing on feature interaction and those centering on user behavior modeling.
 The first category is the feature interaction-based methods, evolving from foundational works such as POLY2 (Chang et al., 2010) and Factorization Machines (FM) (Rendle, 2010).
@@
 The key innovation of these models lies in their ability to identify and utilize complex interactions between a multitude of features to enhance the predictive performance of CTR models.
-Besides, user behavior modeling is another core technique for CTR prediction that mines user preferences from historical interaction behaviors meticulously.
-To better extract informative knowledge from user’s behavior sequence, various network structures have been utilized, including Recurrent Neural Networks (RNNs), Convolutional Neural Networks (CNNs), Attention Networks, and Memory Networks.
-GRU4Rec (Hidasi et al., 2015) designs Gated Recurrent Units (GRUs) to capture the preference-evolving relationship while Caser (Tang and Wang, 2018) leverages the horizontal and vertical convolution to model skip behaviors at both union-level and point-level.
+Furthermore, user behavior modeling is another core technique for CTR prediction that mines user preferences from historical interaction behaviors meticulously.
+To better extract informative knowledge from a user’s behavior sequence, various network structures have been utilized, including Recurrent Neural Networks (RNNs), Convolutional Neural Networks (CNNs), Attention Networks, and Memory Networks.
+GRU4Rec (Hidasi et al., 2015) designs Gated Recurrent Units (GRUs) to capture the preference-evolving relationship, while Caser (Tang and Wang, 2018) leverages the horizontal and vertical convolution to model skip behaviors at both the union-level and point-level.
 Moreover, the attention mechanism is the most popular method for modeling item dependencies, and several influential works have been proposed, including SASRec (Kang and McAuley, 2018), DIN (Zhou et al., 2018), DIEN (Zhou et al., 2019), and BERT4Rec (Sun et al., 2019).
 Among these, DIN and DIEN leverage the target-attention network to identify important historical items, while SASRec and BERT4Rec apply the self-attention network with Transformer architecture to excavate behavior dependencies.
-Besides, memory-based methods (Jiang et al., 2020) are also proposed to store user behavior representations in a read-write manner.
+Additionally, memory-based methods (Jiang et al., 2020) are also proposed to store user behavior representations in a read-write manner.
 2.2. Retrieval-Augmented Recommendation
-To further enhance the performance of CTR prediction, the retrieval-augmented recommendation is proposed, where the most relevant information is retrieved from the historical data.
-Specifically, UBR4CTR (Qin et al., 2020) and SIM (Ren et al., 2019) are designed to retrieve beneficial behaviors from the user’s historical extremely long behavior sequence, where UBR4CTR deploys the search engine method while SIM uses the hard search and soft search approaches.
-To make the search procedure end-to-end, ETA (Chen et al., 2022) is proposed by leveraging the SimHash algorithm to map the user behavior into a low-dimensional space, hence achieving learnable retrieval.
-Moreover, recent works further extend the retrieval-augmented recommendation from item-level retrieval to sample-level retrieval by retrieving informative samples.
-RIM (Qin et al., 2021) is the first to deploy this method that leverages the search engine to retrieve several relevant samples from the search pool and performs neighbor aggregation.
-PET (Du et al., 2022) and DERT (Zheng et al., 2023) have respectively made improvements in the interaction and retrieval mechanism of neighboring samples.
+To further enhance the performance of CTR prediction, retrieval-augmented recommendation is proposed, where the most relevant information is retrieved from historical data.
+Specifically, UBR4CTR (Qin et al., 2020) and SIM (Ren et al., 2019) are designed to retrieve beneficial behaviors from the user’s extremely long historical behavior sequence.
+UBR4CTR deploys the search engine method, while SIM uses hard search and soft search approaches.
+To make the search procedure end-to-end, ETA (Chen et al., 2022) is proposed by leveraging the SimHash algorithm to map user behavior into a low-dimensional space, hence achieving learnable retrieval.
+Moreover, recent works further extend retrieval-augmented recommendation from item-level retrieval to sample-level retrieval by retrieving informative samples.
+RIM (Qin et al., 2021) is the first to deploy this method, leveraging the search engine to retrieve several relevant samples from the search pool and perform neighbor aggregation.
+PET (Du et al., 2022) and DERT (Zheng et al., 2023) have respectively made improvements in the interaction and retrieval mechanisms of neighboring samples.
 PET constructs a hypergraph over the retrieved data samples and performs message propagation to improve the target data representations for final CTR prediction.
@@
 4.1. Overview of ROK
-Despite the superior performance, sample-level retrieval-based methods suffer from the inference inefficiency problem due to the time-consuming online retrieval process.
+Despite their superior performance, sample-level retrieval-based methods suffer from challenges, including inference inefficiency and high resource consumption, due to the time-consuming online retrieval process.
 To this end, we propose a novel Retrieval-Oriented Knowledge (ROK) framework, where a knowledge base is built to imitate the aggregated retrieval knowledge $\mathbf{r}_{t}$.
 As shown in Figure 2, ROK consists of two stages: (1) Retrieval-Oriented Knowledge Construction, and (2) Knowledge Utilization.
-In the first stage of ROK–retrieval-oriented knowledge construction–we pre-train a sample-level retrieval-based model [^2] In this work, we employed RIM(Qin et al., 2021) which is one of the foundational works in sample-level retrieval-based methods.
-The pre-training process exactly follows the original work.
-Importantly, our proposed method extends its applicability to encompass all sample-level retrieval-based methods., followed by designing a delicate knowledge base with a retrieval-oriented embedding layer and a knowledge encoder.
+In the first stage of ROK–retrieval-oriented knowledge construction–we pre-train a sample-level retrieval-based model [^3] In this work, as our research primarily focuses on inference efficiency, we choose a simple yet effective option that provides sufficiently good results: RIM (Qin et al., 2021), which is one of the foundational works in sample-level retrieval-based methods.
+The pre-training process strictly follows the original work (Qin et al., 2021).
+It is important to note that our proposed method can be generalized to all sample-level retrieval-based methods methods, extending its applicability and potential impact., followed by designing a delicate knowledge base with a retrieval-oriented embedding layer and a knowledge encoder.
 This is designed to imitate the retrieval & aggregation results (i.e., $\mathbf{r}_{t}$) of the pre-trained model by directly generating the final aggregated representation $\mathbf{z}_{t}^{x}$ through the knowledge encoder.
@@
 4.2. Structure Design of Knowledge Base
-For sample-level retrieval-based methods (Qin et al., 2021; Du et al., 2022), the overhead of inference inefficiency is mainly caused by the retrieval & aggregation operations, which are uncacheable and unavoidable during the online inference if a brand-new data sample comes.
-To this end, we propose a novel decomposition-reconstruction paradigm, where the aggregated representation $\mathbf{r}_{t}$ is first decomposed into feature-level embeddings, and then reconstructed via a neural encoder.
+For sample-level retrieval-based methods (Qin et al., 2021; Du et al., 2022), the overhead of inference inefficiency and high resource consumption is mainly caused by the retrieval & aggregation operations, which are uncacheable and unavoidable during the online inference if a brand-new data sample comes.
+To this end, we manage to distill knowledge from non-parametric retrieval process using a parametric approach.
+To be specific, we propose a novel decomposition-reconstruction paradigm, where the aggregated representation $\mathbf{r}_{t}$ is first decomposed into feature-level embeddings, and then reconstructed via a neural encoder.
 As illustrated in  3, our proposed knowledge base comprises two key modules: (1) retrieval-oriented embedding layer $f$, and (2) knowledge encoder $g$.
@@
 For simplicity, we adopt multi-layer perceptron (MLP) in this paper.
-This approach reduces the time complexity of sample-level retrieval from $\mathcal{O}(N\log\left(N\right))$, where $N$ is the search pool size to $\mathcal{O}(1)$.
+This approach reduces the time complexity of sample-level retrieval to $\mathcal{O}(1)$.
 Since we introduce the retrieval-oriented embedding layer $f$, we will cut down the embedding size of $f$ from $d$ to $d/2$ for fair comparison in space complexity.
@@
 (3) $\mathcal{L}_{imit}=\mbox{MSE}\left(\mathbf{z}^{x}_{t},\mathbf{r}_{t}\right).$
-This knowledge imitation approach serves as a knowledge distillation process, where we extract and inject the retrieval-enhanced knowledge from the pre-trained model into the learnable knowledge base, enabling the implicit sample-level retrieval capability.
+This knowledge imitation approach acts as a knowledge distillation process, where we distill knowledge from a non-parametric retrieval process using a parametric approach.
+By extracting and injecting the retrieval-oriented knowledge from the pre-trained model into the learnable knowledge base, we enable implicit sample-level retrieval capability.
 4.3.2. Contrastive Regularization
@@
 To prevent collapse in the absence of negative samples (Zhang et al., 2022), a projector $h$ is employed to generate the projected representation $\mathbf{p}^{x}_{t}$, $\mathbf{p}^{s}_{t}$ from $\mathbf{z}^{x}_{t}$, $\mathbf{z}^{s}_{t}$.
-(4) $\vspace*{-0.5\baselineskip}\mathbf{p}^{x}_{t}$ $\triangleq h\left(\mathbf{z}^{x}_{t}\right)\triangleq h\left(g\left(f\left(x_{t}\right)\right)\right),$
+(4) $\mathbf{p}^{x}_{t}$ $\triangleq h\left(\mathbf{z}^{x}_{t}\right)\triangleq h\left(g\left(f\left(x_{t}\right)\right)\right),$
 (5) $\mathbf{p}^{s}_{t}$ $\triangleq h\left(\mathbf{z}^{s}_{t}\right)\triangleq h\left(g\left(f\left(s_{t}\right)\right)\right).\vspace*{-0.5\baselineskip}$
 The cosine similarity between projected representation $\mathbf{p}^{x}_{t}$ and reconstructed representations $\mathbf{z}^{s}_{t}$ is defined as:
-(6) $\mathcal{D}\left(\mathbf{p}^{x}_{t},\mathbf{z}^{s}_{t}\right)=\frac{\mathbf{p}^{x}_{t}}{\left\|\mathbf{p}^{x}_{t}\right\|_{2}}\cdot\frac{\mathbf{z}^{s}_{t}}{\left\|\mathbf{z}^{s}_{t}\right\|_{2}},\vspace*{-0.5\baselineskip}$
+(6) $\mathcal{D}\left(\mathbf{p}^{x}_{t},\mathbf{z}^{s}_{t}\right)=\frac{\mathbf{p}^{x}_{t}}{\left\|\mathbf{p}^{x}_{t}\right\|_{2}}\cdot\frac{\mathbf{z}^{s}_{t}}{\left\|\mathbf{z}^{s}_{t}\right\|_{2}},$
 where $\left\|\cdot\right\|_{2}$ is the $l_{2}$ norm.
@@
 To avoid spatial bias, motivated by the Jensen–Shannon (JS) divergence, a symmetric loss is adopted and the final contrastive regularization loss is:
-(7) $\mathcal{L}_{contra}=-\left(\frac{1}{2}\mathcal{D}\left(\mathbf{p}^{x}_{t},\mathbf{z}^{s}_{t}\right)+\frac{1}{2}\mathcal{D}\left(\mathbf{p}^{s}_{t},\mathbf{z}^{x}_{t}\right)\right).\vspace*{-0.5\baselineskip}$
+(7) $\mathcal{L}_{contra}=-\left(\frac{1}{2}\mathcal{D}\left(\mathbf{p}^{x}_{t},\mathbf{z}^{s}_{t}\right)+\frac{1}{2}\mathcal{D}\left(\mathbf{p}^{s}_{t},\mathbf{z}^{x}_{t}\right)\right).$
 Hence, The overall objective for ROK in the retrieval-oriented knowledge construction stage is:
-(8) $\mathcal{L}=(1-\alpha)\cdot\mathcal{L}_{imit}+\alpha\cdot\mathcal{L}_{contra},\vspace*{-0.5\baselineskip}$
+(8) $\mathcal{L}=(1-\alpha)\cdot\mathcal{L}_{imit}+\alpha\cdot\mathcal{L}_{contra},$
 where $\alpha$ adjusts the ratio of the two loss terms. After the retrieval-oriented knowledge construction stage, the knowledge base will be retained for the follow-up knowledge utilization stage.
@@
   • Feature-wise retrieval-enhanced method leverages the retrieval-oriented embedding layer $f$ to obtain the retrieval-enhanced feature-wise embeddings for the target sample.
-Specifically, this method integrates the retrieval-oriented embedding layer $f$ with the original learnable feature embedding of the backbone model to form the final embedding layer.
-333 The embedding layer accounts for a large proportion of the total number of parameters in CTR models.
+Specifically, this method integrates the retrieval-oriented embedding layer $f$ with the original learnable feature embedding of the backbone model to form the final embedding layer444 The embedding layer accounts for a large proportion of the total number of parameters in CTR models.
 For a fair comparison in model parameters, the embedding size of both the retrieval-oriented embedding layer and the original learnable feature embedding of the backbone model is set to $d/2$ compared with other models.
@@
 These unique representations are then concatenated with the feature embeddings and the combined results are fed into the interaction and MLP (Multi-Layer Perceptron) layers.
-    $\left\{\begin{array}[]{ll}\mathbf{z}_{t}^{x}&=g(f(x_{t}))\\ \hat{y}_{t}&=\text{backbone}(\phi(x_{t})\oplus\mathbf{z}_{t}^{x})\end{array},\right.$
+    (11) $\left\{\begin{array}[]{ll}\mathbf{z}_{t}^{x}&=g(f(x_{t}))\\ \hat{y}_{t}&=\text{backbone}(\phi(x_{t})\oplus\mathbf{z}_{t}^{x})\end{array},\right.$
     where $\phi$ is utilized to extract the unique representations of the backbone model.
@@
 To gain more insights into ROK, we endeavor to address the following research questions (RQs) in this section.
-  • RQ1: How does ROK’s performance compare to that of traditional CTR models and retrieval-based methods?
+  • RQ1: How does ROK’s performance compare to that of traditional CTR models and the teacher model?
   • RQ2: How compatible is ROK with other backbone models?
   • RQ3: How does the knowledge captured by ROK contribute to improving performance?
-  • RQ4: How does the chosen update strategy and positive sample selecting strategy affect ROK’s performance?
-  • RQ5: How do the specified hyperparameter influence ROK’s performance?
+  • RQ4: How do the chosen update strategy and positive sample selecting strategy affect ROK’s performance?
+  • RQ5: How does the specified hyperparameter influence ROK’s performance?
 5.1. Experimental Settings
 5.1.1. Datasets.
-The evaluations were conducted on three widely-recognized public datasets: Tmall, Taobao, and Alipay.
+The evaluations are conducted on three widely-recognized public datasets: Tmall, Taobao, and Alipay.
 Comprehensive statistics for these datasets are presented in Table 1.
-Specifically, we tallied the counts of users, items, samples, fields, and categories for each dataset.
-It’s noteworthy to mention that the number of features refers to the count of unique feature values.
+To be specific, we tally the counts of users, items, samples, fields, and categories for each dataset.
+It’s worth noting that the number of features refers to the count of unique feature values.
 Table 1. The dataset statistics.
@@
 Alipay 498,308 2,200,291 35,179,371 6 3,327,205
-Following RIM (Qin et al., 2021), we organized the data such that the oldest instances constituted the search pool, the most recent instances comprised the test set, and those in between were assigned to the training set.
+Following RIM (Qin et al., 2021), we organize the data such that the oldest instances constitute the search pool, the most recent instances comprised the test set, and the instances in between are assigned to the training set.
 The retrieval-based methods (Qin et al., 2021, 2020) retrieve neighboring samples from the search pool.
-For user behavior modeling methods (Zhou et al., 2018, 2019) involved (e.g., DIEN), the sequential features (e.g., user behavior) are also generated from the search pool.
+For user behavior modeling methods (Zhou et al., 2018, 2019) (e.g., DIEN), the sequential features (e.g., user behavior) are also generated from the search pool.
 5.1.2. Evaluation Metrics.
-To measure the performance, we utilize the commonly adopted metrics AUC and log-loss, which reflect pairwise ranking performance and point-wise likelihood, respectively.
+To measure performance, we utilize the commonly adopted metrics AUC and log-loss, which reflect pairwise ranking performance and point-wise likelihood, respectively.
 A significance test contrasting the two top-performing methods for each metric is undertaken, with results of significance denoted by an asterisk (*).
-We follow (Zhou et al., 2018) to introduce Rel. Impr. metric to quantify the relative improvement of models, which is defined as below:
+We follow previous work (Zhou et al., 2018) to introduce Rel. Impr. metric to quantify the relative improvement of models, which is defined as below:
 […]
@@
 For baselines, we compare ROK with a mid-tier performance backbone model against traditional models and retrieval-based models.
-For traditional CTR models, GBDT (Chen et al., 1996) is a frequently utilized tree-based model, and DeepFM (Guo et al., 2017) is a popular deep-learning model that focuses on feature interactions.
-HPMN (Ren et al., 2019) and MIMN (Pi et al., 2019) utilize memory network architectures, while DIN (Zhou et al., 2018) and DIEN (Zhou et al., 2019) are attention-based CTR models designed to pinpoint user interests.
+For traditional CTR models, GBDT (Chen et al., 1996) is a frequently used tree-based model, and DeepFM (Guo et al., 2017) is a popular deep-learning model that focuses on feature interactions.
+HPMN (Ren et al., 2019) and MIMN (Pi et al., 2019) utilize memory network architectures, while DIN (Zhou et al., 2018) and DIEN (Zhou et al., 2019) are attention-based CTR models designed for pinpointing user interests.
 Additionally, FATE (Wu et al., 2021) is a model for learning representations of tabular data which facilitates interactions among samples within a minibatch.
 In item-level retrieval-based models, SIM (Pi et al., 2020) and UBR (Qin et al., 2020) extract pertinent user behaviors from a comprehensive set of user-generated data.
-As for the sample-level retrieval-based model, RIM (Qin et al., 2021) retrieves relevant data instances based on raw input features, and DERT (Zheng et al., 2023) retrieves dense representations.
-To ensure a fair comparison, we adopted the same hyper-parameter settings for the baselines and the ROK backbone during the knowledge utilization phase as those used in prior work (Qin et al., 2021).
-Regarding backbone models, we intentionally select average-performing common models like DIN and DIEN rather than top-tier ones like UBR, which will further improve the performance of ROK.
-By default, we choose DIEN (Zhou et al., 2019) for Tmall and Taobao datasets and DIN (Zhou et al., 2018) for the Alipay dataset.
-The substantial improvement observed in 5.2 when these backbone models are combined with ROK underscores the efficacy of ROK.
+Regarding the sample-level retrieval-based model, RIM (Qin et al., 2021) retrieves relevant data instances based on raw input features.
+We do not include DERT (Zheng et al., 2023) in our comparison because it is a special case of RIM that introduces a new dense retrieval setting.
+Moreover, our primary goal is to develop a method that eliminates the need for retrieval during inference, thereby improving efficiency and enabling real-world deployment.
+Therefore, we focus more on comparing the performance of ROK with the teacher model i.e., RIM (Qin et al., 2021) to determine whether ROK would cause a significant performance degradation and whether it could enhance the backbone model.
 5.2. Overall Performance Comparison: RQ1
-In this section, we compare the overall experimental results in Table 2, from which we have the following observations.
-First, ROK remarkably boosts the backbone model with AUC improved by 10.08%, 34.96%, and 16.85% on the three datasets respectively, demonstrating the superior performance.
-Second, retrieval-based methods (Qin et al., 2020, 2021) significantly outperform user behavior modeling methods (Zhou et al., 2018, 2019), benefiting from their superior knowledge retrieval ability.
-Moreover, sample-level retrieval-based methods (Qin et al., 2021) perform better than item-level retrieval-based methods (Pi et al., 2020; Qin et al., 2020) due to more comprehensive retrieval.
-Finally, ROK markedly outperforms the item-level retrieval-based methods such as SIM (Pi et al., 2020) and UBR (Qin et al., 2020), and achieves performance surpassing that of the sample-level retrieval-based methods like RIM (Qin et al., 2021).
-Although ROK is only based on knowledge distillation and contrastive learning, it not only surpasses the teacher model but also comes very close to recent results in terms of performance.
-Notably, in our methodology, we deliberately opted for models of average performance, such as DIN and DIEN, over elite models like UBR (Qin et al., 2020), with the anticipation that the latter would enhance the efficacy of ROK.
-Concurrently, it is pertinent to note that the teacher model employed within ROK is RIM, which DERT outperforms.
-Nonetheless, the adoption of a superior teacher model is poised to markedly augment the performance of ROK.
-Despite this, ROK has successfully enabled user behavior modeling methods to outperform both item-level retrieval-based methods and sample-level retrieval-based methods.
-This indicates that the neural knowledge model successfully serves as a compact surrogate for the retrieval pool.
-More importantly, the sample-level retrieval-based methods, due to their prolonged inference time, were once deemed entirely infeasible for industrial application, which will further be elaborated in Section 6.2.
-However, with the introduction of ROK, this barrier has been overcome.
-ROK not only mitigates the challenges but revolutionizes the deployment prospects, rendering sample-level retrieval-based methods feasible industry applications.
+In this section, we compare the overall experimental results presented in Table 2 and make the following observations.
+First, ROK significantly enhances the performance of the backbone model, with AUC improvements of 10.08%, 34.96%, and 16.85% on the three datasets, respectively, demonstrating its superior performance.
+Second, retrieval-based methods (Qin et al., 2020, 2021) considerably outperform user behavior modeling methods (Zhou et al., 2018, 2019), benefiting from their advanced knowledge retrieval capabilities.
+Furthermore, sample-level retrieval-based methods (Qin et al., 2021) exhibit better performance than item-level retrieval-based methods (Pi et al., 2020; Qin et al., 2020) due to more comprehensive retrieval strategies.
+Finally, ROK notably surpasses the performance of item-level retrieval-based methods methods, such as SIM (Pi et al., 2020) and UBR (Qin et al., 2020), and even outperforms our teacher model, RIM (Qin et al., 2021), which belongs to the sample-level retrieval-based methods.
+Our primary goal was to develop a method that eliminates the need for retrieval during inference, thereby improving efficiency and enabling real-world deployment.
+Initially, we expected that our approach would merely approximate the performance of the teacher model while preserving extraordinary inference efficiency, given that it is based on knowledge distillation.
+However, through the introduction of contrastive regularization, our method not only matches but exceeds the performance of the teacher model.
+This demonstrates the strong potential of our approach for real-world applications, as it can effectively learn and utilize the knowledge in the search pool without the need to perform real retrieval during inference.
+Our results highlight the effectiveness of knowledge distillation combined with contrastive regularization in capturing and leveraging the information from the search pool, enabling efficient and high-performing models suitable for deployment in practical scenarios.
 Table 2.
@@
 Evaluation metrics include AUC and log-loss (LL).
-Among the models excluding recent work, the best results are highlighted in bold, while the runner-ups are underlined.
-The gray-colored bold text indicates recent results outperform all other models.
+The best results are highlighted in bold, while the runner-ups are underlined.
 “Rel.
@@
 RIM (Teacher model) 0.9151 0.3697 8.13% 0.8567 0.4546 42.34% 0.8005 0.5736 13.52%
-[…]
 Our Model
@@
 5.3. Compatibility Analysis: RQ2
-In this section, we commence by contrasting the compatibility of our novel approach, ROK, with conventional CTR models, with the findings delineated in Table 3.
+In this section, we begin by contrasting the compatibility of our novel approach, ROK, with conventional CTR models, as delineated in Table 3.
 Upon examining the data, two observations arise: (i) The application of ROK markedly amplifies the performance of the backbone models.
-To illustrate, by integrating ROK with DeepFM, DIN, and DIEN, there’s a notable enhancement in performance metrics, with an average AUC increase of 4.30%, 8.01%, and 5.59%, coupled with an average decrease in log-loss by 11.80%, 8.51%, and 8.22% across three datasets, respectively.
-This highlights the significant benefits of incorporating the retrieval-oriented knowledge from ROK into the backbone models.
-(ii) Exhibiting model-agnosticism, ROK seamlessly integrates with a diverse array of backbone CTR models, irrespective of whether they focus on feature interaction or behavior modeling.
+To illustrate, by integrating ROK with DeepFM (Guo et al., 2017), DIN (Zhou et al., 2018), and DIEN (Zhou et al., 2019), there is a notable enhancement in performance metrics, with an average AUC increase of 4.30%, 8.01%, and 5.59%, coupled with an average decrease in log-loss by 11.80%, 8.51%, and 8.22% across three datasets, respectively.
+This highlights the significant benefits of incorporating the retrieval-oriented knowledge from ROK into the backbone models and indicates that the neural knowledge model successfully serves as a compact surrogate for the search pool.
+(ii) ROK, exhibiting model-agnosticism, seamlessly integrates with a diverse array of backbone CTR models, irrespective of whether they focus on feature interaction or behavior modeling.
 This underscores that the knowledge garnered by ROK is beneficial across diverse models, highlighting the compatibility of ROK.
@@
 In this section, we delve into the comparative analysis of the knowledge quality acquired by ROK and RIM.
-For a detailed assessment, we formulated two simplest variants (LR and MLP) for each model to fully demonstrate the effect of knowledge:
-  • For ROK, we designed the ROK(LR) and ROK(MLP) variants.
-Both of these take as input the concatenation of the original feature embeddings $\mathbf{x}_{t}$, with ROK’s knowledge component $z_{t}^{x}$.
-  • Correspondingly, for RIM, we introduced the RIM(LR) and RIM(MLP) variants.
+For a detailed assessment, we formulate two simple variants (LR and MLP) for each model to fully demonstrate the effect of knowledge:
+  • For ROK, we design the ROK(LR) and ROK(MLP) variants. Both take as input the concatenation of the original feature embeddings $\mathbf{x}_{t}$, with ROK’s knowledge component $\mathbf{z}^{x}_{t}$.
+  • Similarly, for RIM (Qin et al., 2021), we introduced the RIM(LR) and RIM(MLP) variants.
 Their input comprises the concatenation of the original feature embeddings, $\mathbf{x}_{t}$, with RIM’s aggregated features and labels, $\mathbf{r}_{t}$.
-The detailed performance comparison of these variants is tabulated in Table 4 and we have two observations.
+The detailed performance comparison of these variants is tabulated in Table 4 and we make two observations.
 First, significantly, the results indicate that the ROK variants surpass the RIM variants in performance across both datasets.
 This distinction in outcomes emphasizes the superior quality of knowledge learned by ROK, especially as these models apply this intrinsic knowledge directly for predictions.
-It’s compelling to note that while ROK leverages the retrieval imitation strategy (detailed in Section 4.3.1) to imitate the aggregated features and labels of RIM, it extracts richer and more profound knowledge from the data.
-We are inclined to believe that the contrastive regularization methodology is pivotal in driving this enhanced performance for ROK.
-Second, ROK(MLP) in Tmall even outperforms DeepFM and some user behavior modeling methods, which shows another way to improve the pre-ranking model effect.
-To delve deeper into the inherent knowledge of ROK and explain the superiority of ROK over RIM, we employed t-SNE visualization (Van der Maaten and Hinton, 2008).
-Figure 4 and Figure 7 depict the knowledge distribution patterns of both ROK and RIM across the Alipay and Tmall datasets.
+It is compelling to note that while ROK leverages the retrieval imitation strategy (detailed in Section 4.3.1) to imitate the aggregated features and labels of RIM, it extracts richer and more profound knowledge from the data.
+We believe that the contrastive regularization methodology is pivotal in driving this enhanced performance for ROK.
+Second, ROK(MLP) in Tmall even outperforms DeepFM (Guo et al., 2017) and some user behavior modeling methods, which shows another way to improve the pre-ranking model effect.
+To delve deeper into the inherent knowledge of ROK and explain its superiority over RIM, we employed t-SNE visualization (Van der Maaten and Hinton, 2008).
+Figure 4 illustrates the knowledge distribution patterns of both ROK and RIM in the Alipay dataset.
 For this visualization, we randomly selected $10,000$ samples from both the training and testing sets for each model.
 We have two observations.
-First, a discerning observation from Figure 4 reveals that ROK’s knowledge distribution exhibits a more pronounced clustering effect across all datasets.
-This enhanced clustering can be attributed to contrastive regularization.
-For comparison, RIM’s retrieved neighboring samples may include some incidental noise that correlates with the quantity of these samples.
-Besides, the aggregation mechanism of RIM only captures the global features of the neighboring samples.
-However, the contrastive regularization in ROK emphasizes the local features of neighboring samples and mitigates the noise of neighboring samples by selectively considering the most relevant sample, as shown in section 4.3.2.
+First, ROK’s knowledge distribution exhibits a more pronounced clustering effect compared to RIM.
+This enhanced clustering can be attributed to the contrastive regularization employed by ROK.
+The contrastive regularization emphasizes the local features of neighboring samples and mitigates the noise by selectively considering the most relevant sample, as shown in section 4.3.2.
 Specifically, representations of analogous samples are pulled closer, while those of dissimilar samples are pushed apart, leading to a more distinct and meaningful clustering.
-Secondly, a significant concern when using neural networks as a knowledge base in our case relates to their generalization abilities.
-This is because the training process is guided by the retrieved examples of sample-level retrieval-based methods, while the testing phase requires the model to apply its knowledge independently.
-Nonetheless, the visualization of ROK’s representations reveals a high degree of similarity between the training and test datasets, showing the good generalization of the decomposition-reconstruction paradigm.
+In contrast, RIM’s retrieved neighboring samples may include some incidental noise that correlates with the quantity of these samples, and its aggregation mechanism only captures the global features of the neighboring samples.
+Second, the visualization of ROK’s representations reveals a high degree of similarity between the training and test datasets, demonstrating the good generalization ability of the decomposition-reconstruction paradigm.
+This is significant because a major concern when using neural networks as a knowledge base is their generalization ability, as the training process is guided by the retrieved examples of sample-level retrieval-based methods, while the testing phase requires the model to apply its knowledge independently.
 Table 4. Comparison of model variants. “Rel. Impr.” means relative AUC improvement of ROK’s variants against baselines and RIM’s variants. Improvements are statistically significant with $p<0.01$.
@@
 ROK(MLP) 0.8643 - 0.6923 -
-Refer to captionRefer to caption
 ((a)) Alipay (Train)
-Refer to captionRefer to caption
 ((b)) Alipay (Test)
@@
 In this section, we undertake comprehensive ablation experiments for ROK.
-We start by examining the parameter update strategy employed for the knowledge base during the knowledge utilization phase.
+We begin by examining the parameter update strategy employed for the knowledge base during the knowledge utilization phase.
 The strategies under consideration include: fixing the knowledge base (Fix), updating only the upper knowledge encoder $g$ (Upd $g$), and updating both the knowledge encoder and the retrieval-oriented embedding layer $f$ (Upd $f+g$).
-Furthermore, we consider the positive sample selection strategy mentioned in Section 4.3.2.
-Specifically, we select the most related neighboring sample by default.
-Now, we compare the performance of randomly selecting a sample from the retrieved samples (random).
-Results obtained on the Tmall and Alipay datasets are depicted in Figure 5.
-We have two observations: First, we observe that randomly selecting a sample as the positive sample diminishes performance.
-This suggests that even within the narrowly retrieved set of samples, noise is still present.
-Second, a key observation from these results is that the strategy of fixing the knowledge base (Fix) yields the best performance.
+Furthermore, we consider the positive sample selection strategy described in Section 4.3.2.
+Specifically, we select the most related neighboring sample by default and compare its performance to randomly selecting a sample from the retrieved samples (random).
+Figure 5 depicts the results obtained on the Tmall, Taobao and Alipay datasets.
+We have two observations: First, randomly selecting a sample as the positive sample diminishes performance, suggesting that even within the small retrieved set of samples, noise is still present.
+Second, a key observation is that the strategy of fixing the knowledge base (Fix) yields the best performance.
 This can be attributed to the primary goal of ROK: to derive superior representations.
@@
 This characteristic is highly desirable and aligns well with our expectations, further enhancing the practicality of ROK.
-This indicates that once ROK has been fully trained, it can function as an independent module.
-This eliminates the need for subsequent modifications, reinforcing its role as an authentic knowledge base.
+It indicates that once ROK has been fully trained, it can function as an independent module, eliminating the need for subsequent modifications and reinforcing its role as an authentic knowledge base.
 ((a)) Tmall
-((b)) Alipay
+((b)) Taobao
+((c)) Alipay
 Figure 5. Comparison of AUC scores for different update strategies.
 5.5.2. Hyperparameter Study: RQ5.
-In this section, we delve into the study of the hyperparameter $\alpha$ pertaining to the balance between the contrastive regularization loss and retrieval imitation loss, represented as $\alpha$.
-As illustrated in Figure 6, for both datasets, when $\alpha$ assumes values of either 00 or $1$—representing the extreme scenarios—the performance of ROK deteriorates.
+In this section, we delve into the study of hyperparameter $\alpha$ pertaining to the balance between the contrastive regularization loss and retrieval imitation loss, represented as $\alpha$.
+As illustrated in Figure 6, for both datasets, when $\alpha$ equals either 00 or $1$—representing the extreme scenarios—the performance of ROK deteriorates.
 This observation underscores the significance of both retrieval imitation loss and contrastive regularization loss to the model’s efficacy.
-((a)) Tmall
+((a)) Taobao
 ((b)) Alipay
@@
 6. Industry Application
-In this section, we will demonstrate the superior performance of ROK in training and inference efficiency.
 6.1. Training Efficiency
-In evaluating model efficiency, it’s vital to look at both inference speed and training time.
-Training ROK involves pre-training a sample-level retrieval-based model, creating retrieval-oriented knowledge, and applying this knowledge.
-Despite these steps, the extra training time is reasonable compared to a standalone model.
+Training ROK involves pre-training a sample-level retrieval-based model, creating retrieval-oriented knowledge, and applying this knowledge to backbone CTR models.
+Despite these steps, the extra training time is reasonable compared to training a standalone model.
 The training duration is detailed in Table 5.
-Experiments were run on an AMD EPYC 7T83 with an RTX 4090, using hyperparameters—learning rate, batch size, weight decay—optimized via grid search as per the RIM study.
-We also implemented early stopping after 3 steps to enhance efficiency.
 Two observations stand out.
@@
 However, this is mitigated by using large bulk sizes and caching, which decreases retrieval times.
-For further efficiency, extending the knowledge base update frequency beyond the backbone model’s could be beneficial.
+For further efficiency, extending the knowledge base update frequency beyond that of the backbone model could be beneficial.
 Table 5.
@@
 6.2. Inference Efficiency
-In the real world of online services, from advertising platforms to recommendation systems, the significance of inference efficiency cannot be ignored.
-The introduction of the knowledge base significantly amplifies the efficiency in sample-level retrieval, trimming down the time complexity from $\mathcal{O}(N\log\left(N\right))$, where $N$ is the search pool size to a constant time $\mathcal{O}(1)$ through bypassing conventional retrieval & aggregation mechanisms.
-Within this context, we thoroughly assess the inference efficiency of various backbone models—DeepFM (Guo et al., 2017), DIEN (Zhou et al., 2019), and DIN (Zhou et al., 2018)—when seamlessly integrated with ROK and juxtaposed against retrieval-based methods such as UBR (Qin et al., 2020), RIM (Qin et al., 2021), and DERT (Zheng et al., 2023).
-The comparisons drawn from our analyses, as detailed in Table 6, offer insights.
-Specifically, due to the $\mathcal{O}(N\log\left(N\right))$ complexity of the online retrieval process, the inference time of retrieval-based methods is significantly longer than other baseline models.
-For instance, RIM’s inference time is an overwhelming $1\sim 2$ orders of magnitude longer than DeepFM, DIEN, and DIN.
-This significant delay, despite any performance enhancements, fails to align with the real-time response expectations of RSs.
-These systems necessitate a swift response, making such latencies of retrieval-based methods industrially infeasible.
-Moreover, deploying retrieval-based methods for real-world recommendation systems remains complex and costly due to challenges such as encoding discretized features to dense representations, high resource consumption, and divergent inference processes.
-In contrast, our innovative ROK distinguishes itself by delivering notable improvements over the backbone models while maintaining negligible additional inference time, making sample-level retrieval-based methods feasible in online service deployment.
-As shown in Table 6, ROK consistently improves the AUC of backbone models across all datasets while only slightly increasing the inference speed by 0.12ms to 0.25ms per sample.
-Compared to retrieval-based methods, ROK achieves comparable or even better performance with significantly faster inference speed.
-For example, on the Tmall dataset, DIEN+ROK surpasses the performance of DERT (0.9226 vs.
-0.9200) while being more than 3 times faster in inference (5.51ms vs.
-17.78ms per sample).
-ROK’s design allows it to be retrieval-free and compatible with existing recommendation systems.
-It only needs to perform inference once, avoiding additional data transfer overhead and making it highly efficient.
-This represents a blend of predictive accuracy and rapid inference, making ROK a noteworthy solution for modern online services.
-Table 6. Comparison of AUC and inference speed (ms per sample) across models on Tmall, Taobao, and Alipay.
+Deploying sample-level retrieval-based methods is complex and costly due to several key factors like extreme inefficiency problems during inference, high resource consumption and divergent inference process that requires an inference-retrieval-inference cycle.
+Our method, ROK, introduces a knowledge base that improves retrieval efficiency, reducing time complexity to $\mathcal{O}(1)$.
+We assessed backbone models—DeepFM, DIEN, and DIN—integrated with ROK and compared them to UBR (Qin et al., 2020), sparse sample-level retrieval-based methods, and dense sample-level retrieval-based methods.
+As shown in Table 6, retrieval-based methods have significantly longer inference times, making them impractical for real-time requirements.
+In contrast, ROK improves backbone models with minimal additional inference time, making it suitable for online deployment.
+Table 3 and Table 6 demonstrate that ROK consistently improves the AUC of all backbone models across the tested datasets while only slightly increasing the inference time by 0.12 ms to 0.25 ms per sample, showcasing its ability to enhance performance without significantly impacting speed.
+ROK’s retrieval-free design ensures compatibility with existing systems, requiring only one inference process, thus avoiding data transfer overhead and enhancing efficiency.
+This combination of accuracy and rapid inference makes ROK a superior solution for modern online services.
+Table 6.
+Comparison of inference speed (ms per sample) across models on Tmall, Taobao, and Alipay.
+For UBR, sparse retrieval and dense retrieval, both total inference and retrieval times are presented to highlight the impact of the retrieval process.
 Model Tmall Taobao Alipay
-AUC Inference Speed AUC Inference Speed AUC Inference Speed
-DeepFM 0.8585 1.34 0.6710 1.36 0.6971 1.19
-DeepFM+ROK 0.8946 1.46 0.6946 1.40 0.7361 1.41
-DIEN 0.8839 5.44 0.7506 4.89 0.7485 3.69
-DIEN+ROK 0.9226 5.51 0.8382 5.04 0.7884 3.73
-DIN 0.8796 1.28 0.7433 1.32 0.7647 1.43
-DIN+ROK 0.9171 1.46 0.8091 1.36 0.8093 1.56
-UBR (Retrieval) 0.8975 20.71 (17.67) 0.8169 56.45 (53.32) 0.7952 30.32 (27.31)
-RIM (Retrieval) 0.9151 174.81 (173.27) 0.8567 206.22 (204.78) 0.8005 113.95 (112.37)
-DERT (Retrieval) 0.9200 17.78 (16.19) 0.8647 19.53 (16.97) 0.8087 17.95 (16.31)
+UBR (Retrieval) 20.71 (17.67) 56.45 (53.32) 30.32 (27.31)
+Sparse (Retrieval) 174.81 (173.27) 206.22 (204.78) 113.95 (112.37)
+Dense (Retrieval) 17.78 (16.19) 19.53 (16.97) 17.95 (16.31)
+DeepFM 1.34 1.36 1.19
+DeepFM+ROK 1.46 1.40 1.41
+DIEN 5.44 4.89 3.69
+DIEN+ROK 5.51 5.04 3.73
+DIN 1.28 1.32 1.43
+DIN+ROK 1.46 1.36 1.56
 7. Conclusion
-This paper introduces the ROK framework, addressing the inefficiency challenges inherent in sample-level retrieval-based methods during inference.
-By developing a neural network-based knowledge base, ROK imitates and preserves the essential retrieval representations through a novel decomposition-reconstruction approach, further optimized via knowledge distillation and contrastive learning.
-This allows for the integration of retrieval-oriented knowledge with various CTR models, either at an instance or feature level.
-Through rigorous testing on three public datasets, ROK not only matches the performance of existing retrieval-based CTR models but also demonstrates superior inference efficiency.
-These findings highlight ROK’s potential to significantly improve the efficiency and effectiveness of CTR models.
-Appendix A Appendix
-A.1. T-sne visualization
-Figure 7 depict the knowledge distribution patterns of both ROK and RIM in Tmall dataset.
-Refer to captionRefer to caption
-((a)) Tmall (Train)
-Refer to captionRefer to caption
-((b)) Tmall (Test)
-Figure 7. The t-SNE visualization of knowledge from RIM (on the left in each subfigure) and ROK (on the right in each subfigure) for Tmall.
-A.2. Ablation & Hyper-parameter Study on Taobao
-In this section, we present the ablation and hyper-parameter study conducted on the Taobao dataset. The results are depicted in Figure 8, Figure 9.
-Figure 8. Comparison of AUC scores for different update strategies on Taobao.
-Figure 9.
-Comparison of AUC scores across varying $\alpha$ values on Taobao.
-At $\alpha=0$, only Retrieval Imitation loss is utilized, whereas at $\alpha=1$, only the Contrastive Regularization loss is applied.
+The ROK framework tackles the critical challenges faced by sample-level retrieval-based methods, such as inference inefficiency and high resource consumption.
+By introducing a neural network-based Knowledge Base, ROK effectively distills and stores retrieval-oriented knowledge from pre-trained retrieval-based models.
+This innovative approach enables the seamless integration of retrieval-enhanced representations with various CTR models at both the instance and feature levels, bypassing the need for time-consuming retrieval during inference.
+Extensive experiments on three large-scale datasets demonstrate ROK’s exceptional compatibility and performance, significantly enhancing the performance of various CTR methods.
+Remarkably, the introduction of contrastive regularization allows ROK to surpass the performance of the sample-level retrieval-based teacher model, highlighting its strong potential for real-world applications.
+The ROK framework successfully transforms sample-level retrieval-based methods into practical, efficient, and scalable solutions for industrial deployment.
+By eliminating the need for resource-intensive retrieval during inference and demonstrating the feasibility of distilling knowledge from non-parametric models using a parametric approach, ROK opens up new possibilities for leveraging vast amounts of historical data in CTR prediction.
+8. Acknowledgement
+We thank MindSpore (min, 2020) for the partial support of this work, which is a new deep learning computing framework.